# Magnetic correlations and superconducting pairing near higher-order Van Hove singularities

### v1

@@ Abstract
We explore magnetic correlations and superconducting pairing near higher-order Van Hove singularities in an extended Hubbard model on honecycomb lattice incorporating third-nearest-neighbor hopping $t^{\prime\prime}$. Using quantum Monte Carlo methods, we identify a crossover between ferromagnetic and antiferromagnetic fluctuations near higher-order Van Hove singularities filling, where $t^{\prime\prime}$ enhances ferromagnetic correlations below while suppressing antiferromagnetic fluctuations toward half-filling. At low doping, $f_{n}$-wave pairing dominates, amplified by higher-order Van Hove singularities-induced divergent density of states. Remarkably, despite general suppression of $f_{n}$-wave pairing by increasing next-nearest neighbor hopping $t^{\prime}$ and $t^{\prime\prime}$, a critical $t^{\prime\prime}=0.15$ triggers anomalous enhancement via higher-order Van Hove singularities renormalization at a fix $t^{\prime}$. The nearest-neighbor Coulomb interactions $V$ suppress superconducting correlation, which exhibiting sign-independent suppression proportional to $|V|$. These results highlight the interplay of higher-order Van Hove singularities-driven electronic structure, magnetic fluctuations, and pairing symmetry competition in electron correlated systems.
@@ I Introduction
In the realm of strongly correlated electron systems, the exploration of unconventional superconductivity stands as one of the most captivating frontiers in condensed matter physics[1]. Unlike conventional superconductors, where phonon-mediated pairing gives rise to Bardeen-Cooper-Schrieffer (BCS) superconductivity, unconventional superconductors typically emerge from intricate electronic interactions, yielding exotic pairing symmetries and non-trivial order parameters[2, 1]. Extensive investigations have demonstrated that magnetic fluctuations driven by strong electron correlations play a pivotal role in mediating Cooper pairing, as seen in doped cuprates[3], iron-based superconductors[4], and heavy-fermion systems[5]. Specifically, triplet pairing may arise from ferromagnetic-like spin fluctuations, while singlet pairing with $d$-wave symmetry is often associated with antiferromagnetic spin fluctuations[1, 6]. This dichotomy undergoes the ongoing debate about the competition between different superconducting pairing symmetries and magnetic modes.
@@
In recent years, an especially intriguing scenario has emerged in materials with higher-order van Hove singularities (HOVH), where the electronic DOS diverges in a power-law manner due to the presence of saddle points with higher-order band flattening[16, 7, 17, 18, 19]. Unlike conventional van Hove singularities, which exhibit a logarithmic DOS divergence. HOVH systems can sustain markedly enhanced electronic correlations, amplifying the effects of interactions such as spin fluctuations, charge instabilities, and unconventional superconductivity[20, 21].
@@
Theoretical and experimental evidence is emerging to suggest that strong electron-electron interactions induced by HOVH may give rise to a distinct type of chiral superconductivity, characterized by a unique pairing symmetry that deviates from conventional superconducting mechanisms[22]. These singularities have also been shown to trigger quantum phase transitions, leading to substantial modifications in the topological properties of materials and even the emergence of topological superconductivity[14, 23]. The HOVH not only enhances the electronic anisotropy of materials, instigating a competition between ferromagnetic (FM) and antiferromagnetic (AFM) fluctuations[7], but also promotes Stoner-type magnetism due to the extreme DOS. Moreover, it provides a favorable environment for long-range electron pairing, potentially stabilizing superconducting phases beyond the standard BCS paradigm[1].
@@
Inspired by the intriguing correlated phenomena discovered in graphene-based systems, the Hubbard model on a honeycomb lattice has become a paradigmatic theoretical framework for exploring strong electron correlations in quantum materials. Beyond its Dirac physics at half filling, this prototypical system exhibits VHS in its low-energy spectrum, with the saddle points of the band structure leading to a divergent DOS[24]. Notably, when these saddle points acquire higher-order characteristics (e.g., through strain, twisting, or doping), the singularities evolve into higher-order Van Hove singularities, leading to a more dramatic amplification of the DOS. In such regimes, the enhanced electronic correlations near HOVH create a fertile platform for intertwined orders, including unconventional superconductivity and complex magnetic phases[25].
@@
The Hubbard model on a honeycomb lattice has been intensively studied for decades. At half-filling, quantum Monte Carlo (QMC) studies indicate that when the on-site Coulomb repulsion energy $U$ exceeds the critical value $U_{c}\approx 3.8t$, where $t$ is the nearest-neighbor(NN) hopping, the system undergoes a transition from a semimetal to a Mott insulator, accompanied by the emergence of AFM long-range order[26, 27, 28]. When the system is away from half-filling, electron correlations may induce unconventional superconductivity: near half-filling, superconducting pairing is often dominated by the $d+id$ wave (chiral $d$ wave)[29, 25, 30, 31, 32, 33, 34, 35, 36, 37], which originates from the mediation of AFM spin fluctuations. In the large-doping region ($\langle n\rangle\simeq 0.2$), when the Fermi level is close to the HOVH singularity at the bottom of the energy band induced by the next-nearest-neighbor (NNN) hopping term $t^{\prime}$, the $p+ip$ wave (chiral $p$ wave) triplet pairing becomes dominant[6, 38, 39, 40]. This dominant superconducting pairing is driven by FM fluctuations, and the power-law divergence of the DOS near the HOVH may significantly increase the superconducting critical temperature. However, although the role of HOVH in enhancing electron correlations and superconducting pairing has been preliminarily explored, the reported filling region is rather low, making experimental detection nearly impossible[6].
@@
In honeycomb lattices, the complex interplay between HOVH singularities and correlation-driven instabilities—such as superconductivity and magnetic order—remains a profound mystery. This work expands the existing framework by incorporating the third-nearest-neighbor (TNN) hopping parameter $t^{\prime\prime}$: using the determinant quantum Monte Carlo (DQMC) method, we investigate the impact of $t^{\prime\prime}$ on magnetic correlations and observe a crossover between FM and AFM fluctuations at different electron fillings. By employing constrained-path Monte Carlo method, we focus on electron correlation-driven superconducting pairing near HOVH under various conditions of electron filling $\langle n\rangle$, hopping parameters $t^{\prime}$ and $t^{\prime\prime}$, and nearest-neighbor Coulomb interaction $V$. Our findings reveal that near the HOVH filling, superconducting pairing with $f_{n}$-symmetry becomes dominant, and this pairing may be driven by competing FM/AFM fluctuations.
@@ II Model and Methods
The system we simulated is sketched in Fig.1 (a), and the honeycomb lattice structure consists of two sublattices: the A sublattice, represented by blue circles, and the B sublattice, denoted by red circles. The first Brillouin zone (BZ), depicted in Fig. 1(b), shows high-symmetry points along standard paths, essential for electronic structure analysis.
Figure 1: (a) Sketch of honeycomb lattice with double-27 sites of $N=2\times 3L^{2}$ with $L=3$; (b) the first Brillouin zone and high symmetry points; (c) the form of NN $d+id$ pairing symmetry; (d) the form of NNN $f_{n}$ pairing symmetry.
@@
[…]
$H_{U}$ $=U\sum_{i}n_{i\uparrow}n_{i\downarrow}+V\sum_{\langle i,j\rangle\sigma\sigma^{\prime}}n_{i\sigma}n_{j\sigma^{\prime}}$ (1)
@@
Compared to the pristine honeycomb lattice with only NN hopping $t$, the inclusion of long-range hoppings $t^{\prime}$ and $t^{\prime\prime}$ enhances electron delocalization while intensifying geometric frustration. This dual effect modifies the electronic band structure and shifts the HOVH singularity position. To quantitatively determine the HOVH location, we analyze the non-interacting ($U=0$) band dispersion:
$E_{\pm}(\vec{k})=\pm\mid t\alpha(\vec{k})+t^{\prime\prime}\gamma(\vec{k})\mid-t^{\prime}\beta(\vec{k})$
@@
Here $\alpha(\vec{k})=\sum\limits_{i=1}^{3}e^{-i\vec{k}\cdot\vec{\mathbf{a}}_{i}}$, $\beta(\vec{k})=\sum\limits_{i=1}^{6}e^{-i\vec{k}\cdot\vec{\mathbf{b}}_{i}}$, $\gamma(\vec{k})=\sum\limits_{i=1}^{3}e^{-i\vec{k}\cdot\vec{\mathbf{c}}_{i}}$, with $\vec{\mathbf{a}}_{i}$, $\vec{\mathbf{b}}_{i}$, and $\vec{\mathbf{c}}_{i}$ denote NN, NNN, and TNN displacement vectors, respectively. Expanding $E(\vec{k})$ around the $M$-point up to quadratic order, we identify the HOVH condition when the effective mass tensor becomes singular. Specifically, the critical relation[41]
$t^{\prime\prime}=\frac{t-2t^{\prime}}{4}$ (2)
signals a topological transition at the saddle point, where the conventional VHS evolves into a HOVH through Berry curvature restructuring. This transition is reflected in the change from the black solid line to the red solid line, as demonstrated in Fig. 2.
@@
Figure 2: The DOS as functions of energy with $t=1.00$, $t^{\prime}$, $t^{\prime\prime}=0$ (black solid line) and $t=1.00$, $t^{\prime}=0.10$, $t^{\prime\prime}=0.20$ (red solid line). And the red dotted line shows fillings $\langle n\rangle$ as a function of energy with $t=1.00$, $t^{\prime}=0.10$, and $t^{\prime\prime}=0.20$
@@
In our simulations, we employed a two dimensional honeycomb lattice with periodic boundary conditions. The system size was set to $L=$ 4 and 5, leading to a total of $N=2\times 3L^{2}$ lattice sites. This parameter selection enables us to investigate fundamental physical phenomena within a computationally tractable framework while effectively capturing the essential characteristics of the honeycomb lattice structure. The system is studied using complementary quantum Monte Carlo approaches: DQMC for finite-temperature properties[42, 43, 44, 45] and CPMC for ground-state characterization[46, 47, 48, 33, 48, 47, 29, 49]. The DQMC methodology employs a stochastic evaluation of the partition function through auxiliary field discretization, where the high-dimensional integration is performed via importance sampling with controlled Trotter errors. In contrast, the CPMC approach projects the ground-state wavefunction through imaginary-time evolution within a constrained manifold of Slater determinants, enforced by a trial wavefunction $|\psi_{T}\rangle$ that maintains positive overlaps throughout the random walk process. Extensive benchmarking confirms the constraint-induced systematic error remains below 3% for measured observables, with ground-state properties showing remarkable insensitivity to the specific choice of $|\psi_{T}\rangle$ [42]. Our implementation utilizes closed-shell filling configurations and adopts a free-electron trial wavefunction, $|\psi_{T}\rangle=\prod_{k\leq k_{F}}c_{k}^{\dagger}|0\rangle$, where the Fermi surface corresponds to the non-interacting system. This construction ensures computational efficiency while preserving physical consistency with the metallic phase under investigation, achieving an optimal balance between numerical tractability and systematic error control.
@@
To study magnetic fluctuations, we defined the spin susceptibility in the $z$-direction at zero frequency as follows:
[…]
Here, $\chi_{s}(\Gamma)$ quantifies ferromagnetic correlations, and $\chi_{s}(M)$ quantifies antiferromagnetic correlations.
@@
Within the CPMC framework, we investigate long-range superconducting pairing correlations by analyzing pairing correlation functions for various pairing symmetries. The pairing correlation function is defined as:
$C_{\alpha}(r)=\frac{1}{N}\sum_{i}\langle\Delta_{\alpha}^{\dagger}(\mathbf{r}_{i}+\mathbf{r})\Delta_{\alpha}(\mathbf{r}_{i})\rangle$ (4)
For the NN pairing (e.g., $d+id$ pairing), the pairing operator takes the form[50, 51]:
$\Delta_{d+id}(\mathbf{r}_{i})=\sum_{\delta}\eta_{d+id}(\delta)\left(c_{i\uparrow}c_{i+\delta\downarrow}-c_{i\downarrow}c_{i+\delta\uparrow}\right)$ (5)
For the NNN pairing (e.g., $f_{n}$ pairing), the pairing operator is given by:
$\Delta_{f_{n}}(\mathbf{r}_{i})=\sum_{\delta^{\prime}}\eta_{f_{n}}(\delta^{\prime})\left(c_{i\uparrow}c_{i+\delta^{\prime}\downarrow}+c_{i\downarrow}c_{i+\delta^{\prime}\uparrow}\right)$ (6)
@@
$\eta_{d+id}(\delta)=\begin{cases}1,&\delta=\vec{\mathbf{a}}_{1}\\ e^{i2\pi/3},&\delta=\vec{\mathbf{a}}_{2}\\ e^{i4\pi/3},&\delta=\vec{\mathbf{a}}_{3}\end{cases}$ (7)
$\eta_{f_{n}}(\delta^{\prime})=\sin\left(\frac{\mathbf{k}\cdot\delta^{\prime}}{2}\right),\;\delta^{\prime}=\vec{\mathbf{a}}_{1},\vec{\mathbf{a}}_{2},...\,\vec{\mathbf{a}}_{6}$ (8)
Their pairing symmetry configurations are sketched in Fig.1(c) and Fig.1 (d), respectively.

### v2

@@ Abstract
Higher-order Van Hove singularities in strongly correlated electron systems provide a fertile ground for emergent electronic orders and superconductivity. This study investigates the interplay between magnetic fluctuations and superconducting pairing near higher-order Van Hove singularities on the honeycomb lattice, a paradigmatic platform relevant to graphene. By incorporating third-nearest-neighbor hopping $t^{\prime\prime}$, we uncover a universal crossover: ferromagnetic fluctuations dominate below the higher-order Van Hove filling, while antiferromagnetic fluctuations take over toward half filling. A key finding is that the already dominant $f_{n}$-wave pairing is enhanced in the critical region of this magnetic crossover by the higher-order Van Hove. This enhancement is driven by the synergistic effect of the higher-order Van Hove singularities-induced divergent density of states and the competing magnetic fluctuations. Although increased hopping parameters generally suppress superconducting correlation, we identify a critical $t^{\prime\prime}$ that anomalously enhances pairing via the higher-order Van Hove renormalization. Furthermore, the nearest-neighbor Coulomb interaction suppresses the pairing correlation function in a sign-independent manner. Our results clarify the competitive mechanisms between magnetic fluctuations and unconventional superconductivity in higher-order Van Hove singularities systems, offering a theoretical basis for tailoring quantum phases in graphene-based materials via band engineering.
@@ I Introduction
In the realm of strongly correlated electron systems, the exploration of unconventional superconductivity stands as one of the most captivating frontiers in condensed matter physics[37]. Unlike conventional superconductors, where phonon-mediated pairing gives rise to Bardeen-Cooper-Schrieffer (BCS) superconductivity, unconventional superconductors typically emerge from intricate electronic interactions, yielding exotic pairing symmetries and non-trivial order parameters[37]. Extensive investigations have demonstrated that magnetic fluctuations driven by strong electron correlations play a pivotal role in mediating Cooper pairing, as seen in doped cuprates[2], iron-based superconductors[24], and heavy-fermion systems[43]. Specifically, triplet pairing may arise from ferromagnetic-like spin fluctuations, while singlet pairing with $d$-wave symmetry is often associated with antiferromagnetic spin fluctuations[37, 32]. This dichotomy undergoes the ongoing debate about the competition between different superconducting pairing symmetries and magnetic modes.
@@
In recent years, an especially intriguing scenario has emerged in materials with higher-order Van Hove (HOVH) singularities, where the electronic DOS diverges in a power-law manner due to the presence of saddle points with higher-order band flattening[54, 18, 21, 45, 55]. Unlike conventional VHS, which exhibit a logarithmic DOS divergence, HOVH systems can sustain markedly enhanced electronic correlations, amplifying the effects of interactions such as spin fluctuations, charge instabilities, and unconventional superconductivity[49, 39].
@@
Theoretical and experimental evidence is emerging to suggest that strong electron-electron correlation effects, which are greatly amplified by the extreme DOS induced by HOVH, may give rise to a distinct type of chiral superconductivity, characterized by a unique pairing symmetry that deviates from conventional superconducting mechanisms[34]. These singularities have also been shown to trigger quantum phase transitions, leading to substantial modifications in the topological properties of materials and even the emergence of topological superconductivity[16, 19]. The HOVH not only enhances the electronic anisotropy of materials, instigating a competition between ferromagnetic (FM) and antiferromagnetic (AFM) fluctuations[18], but also promotes Stoner-type magnetism due to the extreme DOS. Moreover, it provides a favorable environment for long-range electron pairing, potentially stabilizing superconducting phases beyond the standard BCS paradigm[37].
@@
Inspired by the intriguing correlated phenomena discovered in graphene-based systems, the Hubbard model on a honeycomb lattice has become a paradigmatic theoretical framework for exploring strong electron correlations in quantum materials. Beyond its Dirac physics at half filling, this prototypical system exhibits VHS in its low-energy spectrum, with the saddle points of the band structure leading to a divergent DOS[30]. Notably, when these saddle points acquire higher-order characteristics (e.g., through strain, twisting, or doping), the singularities evolve into HOVH, leading to a more dramatic amplification of the DOS. In such regimes, the enhanced electronic correlations near HOVH create a fertile platform for intertwined orders, including unconventional superconductivity and complex magnetic phases[23].
While HOVH can be engineered in various lattices such as square, triangular, and kagome[7, 19, 36, 18], the honeycomb lattice is a paradigmatic platform due to its direct relevance to graphene and the clear, controllable pathway to HOVH via further-neighbor hoppings[10]. This work leverages these advantages to systematically unravel the interplay between HOVH, magnetic fluctuations, and pairing symmetry—a complex nexus that remains less explored across lattice geometries.
@@
The Hubbard model on a honeycomb lattice has been intensively studied for decades. At half filling, quantum Monte Carlo (QMC) studies indicate that when the on-site Coulomb repulsion energy $U$ exceeds the critical value $U_{c}\approx 3.8|t|$, where $t$ is the nearest-neighbor (NN) hopping, the system undergoes a transition from a semimetal to a Mott insulator, accompanied by the emergence of AFM long-range order[33, 42, 1]. When the system is away from half filling, electron correlations may induce unconventional superconductivity: near half filling, superconducting pairing is often dominated by the $d+id$ wave (chiral $d$ wave) [31, 23, 15, 14, 11, 12, 17, 51, 28, 50], which originates from the mediation of AFM spin fluctuations. In the large-doping region ($\left\langle n\right\rangle\simeq 0.2$), when the Fermi level is close to the HOVH singularity at the bottom of the energy band induced by the next-nearest-neighbor (NNN) hopping term $t^{\prime}$, the $p+ip$ wave (chiral $p$ wave) triplet pairing becomes dominant[32, 53, 52, 8]. This dominant superconducting pairing is driven by FM fluctuations, and the power-law divergence of the DOS near the HOVH may significantly increase the superconducting critical temperature. However, although the role of HOVH in enhancing electron correlations and superconducting pairing has been preliminarily explored, the reported filling region is rather low, making experimental detection nearly impossible[32].
@@
In honeycomb lattices, the complex interplay between HOVH singularities and correlation-driven instabilities—such as superconductivity and magnetic order—remains a profound mystery. This work expands the existing framework by incorporating the third-nearest-neighbor (TNN) hopping parameter $t^{\prime\prime}$: using the determinant quantum Monte Carlo (DQMC) method, we investigate the impact of $t^{\prime\prime}$ on magnetic correlations and observe a crossover between FM and AFM fluctuations at different electron fillings. By employing constrained-path Monte Carlo (CPMC) method, we focus on electron correlation-driven superconducting pairing near HOVH under various conditions of electron filling $\langle n\rangle$, hopping parameters $t^{\prime}$ and $t^{\prime\prime}$, and nearest-neighbor Coulomb interaction $V$. Our findings reveal that near the HOVH filling, superconducting pairing with $f_{n}$ symmetry becomes dominant, and this pairing may be driven by competing FM/AFM fluctuations.
@@ II Model and Methods
The system we simulated is sketched in Fig. 1(a), and the honeycomb lattice structure consists of two sublattices: the $A$ sublattice, represented by blue circles, and the $B$ sublattice, denoted by red circles. The first Brillouin zone, depicted in Fig. 1(b), shows high-symmetry points along standard paths, essential for electronic structure analysis.
Figure 1: (a) Sketch of the honeycomb lattice with double-27 sites of $N=2\times 3L^{2}$ with $L=3$. (b) The first Brillouin zone and high symmetry points. (c) The form of NN $d+id$ pairing symmetry. (d) The form of NNN $f_{n}$ pairing symmetry.
@@
$-t^{\prime\prime}\sum_{\langle\langle\langle i,j\rangle\rangle\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+\text{H.c.},$
$H_{U}$ $=U\sum_{i}n_{i\uparrow}n_{i\downarrow}+V\sum_{\langle i,j\rangle\sigma\sigma^{\prime}}n_{i\sigma}n_{j\sigma^{\prime}}.$ (1)
@@
In our calculations, the hopping parameters $t^{\prime}$, and $t^{\prime\prime}$ are expressed in units of the nearest-neighbor hopping $t$. For graphene, $t$ is typically $\approx 2.0\ \text{eV}$[32], and the next-nearest-neighbor hopping $t^{\prime}$ is estimated to range from $0.10t$ to $0.30t$ in relevant systems[25, 13]. The critical relation in Eq. (2) dictates the emergence of the HOVH. For a representative $t^{\prime}=0.20t$, it yields a critical $t^{\prime\prime}=0.15t$, which is the value at which we observe the anomalous enhancement of $f_{n}$ wave pairing.
@@
Compared to the pristine honeycomb lattice with only NN hopping $t$, the inclusion of long-range hoppings $t^{\prime}$ and $t^{\prime\prime}$ enhances electron delocalization while intensifying geometric frustration. This dual effect modifies the electronic band structure and shifts the HOVH singularity position. To quantitatively determine the HOVH location, we analyze the noninteracting ($U=0$) band dispersion:
$E_{\pm}(\vec{k})=\pm\mid t\alpha(\vec{k})+t^{\prime\prime}\gamma(\vec{k})\mid-t^{\prime}\beta(\vec{k}).$
@@
Here $\alpha(\vec{k})=\sum\limits_{i=1}^{3}e^{-i\vec{k}\cdot\vec{\mathbf{a}}_{i}}$, $\beta(\vec{k})=\sum\limits_{i=1}^{6}e^{-i\vec{k}\cdot\vec{\mathbf{b}}_{i}}$, and $\gamma(\vec{k})=\sum\limits_{i=1}^{3}e^{-i\vec{k}\cdot\vec{\mathbf{c}}_{i}}$, where $\vec{\mathbf{a}}_{i}$, $\vec{\mathbf{b}}_{i}$, and $\vec{\mathbf{c}}_{i}$ denote NN, NNN, and TNN displacement vectors, respectively. Expanding $E(\vec{k})$ around the $M$ point up to quadratic order, we identify the HOVH condition when the effective mass tensor becomes singular. Specifically, the critical relation[10]
$t^{\prime\prime}=\frac{t-2t^{\prime}}{4}.$ (2)
This critical relation signals a topological transition at the saddle point, where the conventional VHS evolves into a HOVH. To illustrate this transition, we plot in Fig. 2(a) the DOS for two representative cases: without HOVH ($t^{\prime}=t^{\prime\prime}=0$, black solid line) and with HOVH ($t^{\prime}=0.10$, $t^{\prime\prime}=0.20$, red solid line). The transition is reflected in the change from the black solid line to the red solid line. Furthermore, Fig. 2(b) shows the noninteracting band dispersion along the high-symmetry path for different $t^{\prime\prime}$ values at fixed $t=1.0,t^{\prime}=0.20$, providing a direct visualization of the HOVH formation.
@@
Figure 2: (a) The DOS as functions of energy with $t=1.00$, $t^{\prime}=0$, $t^{\prime\prime}=0$ (black solid line) and $t=1.00$, $t^{\prime}=0.10$, $t^{\prime\prime}=0.20$ (red solid line). The red dotted line shows fillings $\langle n\rangle$ as a function of energy with $t=1.00$, $t^{\prime}=0.10$, and $t^{\prime\prime}=0.20$. (b) Noninteracting band dispersions for different $t^{\prime\prime}$ values at fixed $t=1.0,t^{\prime}=0.20$
In our simulations, we employed a two-dimensional honeycomb lattice with periodic boundary conditions, and set the onsite Coulomb interaction to $U=3.0|t|$. This value, which is half the non-interacting bandwidth of the pristine lattice, places the system in a moderately to strongly correlated regime. And the use of $U=3.0|t|$ falls within the well-justified range of U/|t|=13U/|t|=1-3 employed in prior DQMC work by Cheng et al.[9], which explicitly states the reliability of DQMC simulations within this correlated regime. This standard choice ensures strong correlations are present to investigate the interplay between magnetic fluctuations, superconducting pairing, and the HOVH, even with the inclusion of $t^{\prime}$ and $t^{\prime\prime}$. The system size was set to $L=$ 4 and 5, leading to a total of $N=2\times 3L^{2}$ lattice sites. This parameter selection enables us to investigate fundamental physical phenomena within a computationally tractable framework while effectively capturing the essential characteristics of the honeycomb lattice structure.
The choice of QMC methods is particularly suited to this study. Our system resides in an intermediate-to-strong coupling regime ($U=3.0|t|$), where perturbative methods often fail to capture strong electron correlations and fluctuations reliably. QMC provides a numerically exact treatment for such problems. Specifically, DQMC is used for finite-temperature spin susceptibility, as the negative sign problem is not severe, ensuring statistical reliability[41, 3, 44, 38] . Concurrently, CPMC is employed as it effectively circumvents the sign problem, enabling extraction of ground-state superconducting pairing correlations[35, 57, 56, 12, 56, 57, 31, 47]. This combined QMC approach thus offers a rigorous framework for our studies.
@@
The DQMC methodology employs a stochastic evaluation of the partition function through auxiliary field discretization, where the high-dimensional integration is performed via importance sampling with controlled Trotter errors. In contrast, the CPMC approach projects the ground-state wave function through imaginary-time evolution within a constrained manifold of Slater determinants, enforced by a trial wave function $|\psi_{T}\rangle$ that maintains positive overlaps throughout the random walk process. Extensive benchmarking confirms the constraint-induced systematic error remains below 3% for measured observables, with ground-state properties showing remarkable insensitivity to the specific choice of $|\psi_{T}\rangle$ [41]. Our implementation utilizes closed-shell filling configurations and adopts a free-electron trial wave function, $|\psi_{T}\rangle=\prod_{k\leq k_{F}}c_{k}^{\dagger}|0\rangle$, where the Fermi surface corresponds to the non-interacting system. This construction ensures computational efficiency while preserving physical consistency with the metallic phase under investigation, achieving an optimal balance between numerical tractability and systematic error control.
@@
To study magnetic fluctuations, we defined the spin susceptibility in the $z$ direction at zero frequency as follows:
$\chi_{s}(\mathbf{q})=\frac{1}{N}\int_{0}^{\beta}d\tau\sum_{i,j}e^{i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}\langle S_{i}^{z}(\tau)S_{j}^{z}(0)\rangle.$ (3)
Here, $\chi_{s}(\Gamma)$ quantifies FM correlations, and $\chi_{s}(M)$ quantifies AFM correlations.
@@
Within the CPMC framework, we investigate long-range superconducting pairing correlations by analyzing pairing correlation functions for various pairing symmetries. The pairing correlation function is defined as
$C_{\alpha}(r)=\frac{1}{N}\sum_{i}\langle\Delta_{\alpha}^{\dagger}(\mathbf{r}_{i}+\mathbf{r})\Delta_{\alpha}(\mathbf{r}_{i})\rangle.$ (4)
For the NN pairing (e.g., $d+id$ pairing), the pairing operator takes the form[20, 29]
$\Delta_{d+id}(\mathbf{r}_{i})=\sum_{\delta}\eta_{d+id}(\delta)\left(c_{i\uparrow}c_{i+\delta\downarrow}-c_{i\downarrow}c_{i+\delta\uparrow}\right).$ (5)
For the NNN pairing (e.g., $f_{n}$ pairing), the pairing operator is given by
$\Delta_{f_{n}}(\mathbf{r}_{i})=\sum_{\delta^{\prime}}\eta_{f_{n}}(\delta^{\prime})\left(c_{i\uparrow}c_{i+\delta^{\prime}\downarrow}+c_{i\downarrow}c_{i+\delta^{\prime}\uparrow}\right).$ (6)
@@
$\eta_{d+id}(\delta)=\begin{cases}1,&\delta=\vec{\mathbf{a}}_{1}\\ e^{i2\pi/3},&\delta=\vec{\mathbf{a}}_{2},\\ e^{i4\pi/3},&\delta=\vec{\mathbf{a}}_{3}\end{cases}$ (7)
$\eta_{f_{n}}(\delta^{\prime})=\sin\left(\frac{\mathbf{k}\cdot\delta^{\prime}}{2}\right),\;\delta^{\prime}=\vec{\mathbf{a}}_{1},\vec{\mathbf{a}}_{2},...\,\vec{\mathbf{a}}_{6}.$ (8)
Their pairing symmetry configurations are sketched in Figs.1(c) and 1(d), respectively.